OpenStudy (anonymous):

Check answers please! Medal! Solve the equation sqrt11x-5=sqrt9x+7 I got 6

OpenStudy (anonymous):

@neer2890

OpenStudy (neer2890):

\[\sqrt{11x-5}=\sqrt{9x+7}\] is this correct?

OpenStudy (anonymous):

yes

OpenStudy (neer2890):

squaring both sides 11x-5=9x+7 11x-9x=5+7 2x=12 x=6

OpenStudy (anonymous):

Thank you! I have a few more if you dont mind checking. I will post them again

OpenStudy (neer2890):

okay

OpenStudy (neer2890):

in these type of questions apply same rule and you will get the right answer